OpenStudy (anonymous):

if a does not equal b and a#b=10, which of the following could be true? I. a+b is even II. a+B is odd III. b#a=10

OpenStudy (anonymous):

so III. is not right because b#a=10 would equal b, which does not equal a?

OpenStudy (anonymous):

after this question, I have one more problem from this section that I need help on.

OpenStudy (kinggeorge):

Assuming this is still the same operation from the last problem, then III. is not correct because of the following. Let a#b=10. That means that either a=10, or b=10. Now suppose that b#a=10. Since b+a=a+b, we have that a=b. This contradicts our assumption, so III. can't be correct.

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

k, next problem